Solve y2-10y+25=0,using completing the square method
Answer by stanbon(75887)

Solve y2-10y+25=0,using completing the square method
------
y^2 - 10y + 5^2 = -25+5^2
Factor:
(y-5)^2 = 0
----
Solution:
y = 5 with multiplicity two.
=========================
Cheers,
Stan H.